OpenStudy (anonymous):

Emma finished a small bag of her favorite collation that weighed 0.125 pound. Wendy ate 1/7 of a one-pound bag of the snack. Who ate more? Explain.

OpenStudy (pfenn1):

Emma ate 0.125 pounds or 1/8 pound Wendy are 1/7 pound Who ate more?

OpenStudy (pfenn1):

Do you know if 1/7 is larger or smaller than 1/8?

OpenStudy (anonymous):

is 1/8 larger

OpenStudy (pfenn1):

Is 1/2 larger or smaller than 1/4?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

because 4 is larger than 2

OpenStudy (anonymous):

and 1 and 1 are the same

OpenStudy (pfenn1):

Think about a pie that is a kind you really like. Would you rather some one gave you 1/2 of the pie or 1/4 of the pie? Which piece would be bigger? Would you rather split the pie with 1 other friend or with 3 other friends?

OpenStudy (pfenn1):

if you try to find the common denominator for 1/2 and 1/4, it would be 4. If you multiply 1/2 by 2/2 you would get 1/2 = 2/4. So is 2/4 more or less than 1/4?

OpenStudy (pfenn1):

So find the common denominator for 1/7 and 1/8. It would be 7 times 8 or 56. 1/7 * 8/8 = 8/56 1/8 * 7/7 = 7/56 Which is larger?

OpenStudy (anonymous):

1/7

OpenStudy (pfenn1):

Yes!

OpenStudy (anonymous):

ooh

OpenStudy (pfenn1):

So Wendy ate more than Emma!

OpenStudy (anonymous):

so 1/2 is better than 1/4

OpenStudy (pfenn1):

Yes!

OpenStudy (anonymous):

ok

OpenStudy (pfenn1):

I wish I had 1/2 of a pecan pie right now. I'm hungry!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

as soon as you said pie i got hungry

OpenStudy (pfenn1):

LOL! I have to go make dinner. Good luck with the rest of your homework!

OpenStudy (anonymous):

thanks and good bye.

OpenStudy (pfenn1):

you are welcome. bye.